# Low-pass sampling in Model Predictive Path Integral Control

### v2

@@ Abstract
Model Predictive Path Integral (MPPI) control is a widely used sampling-based approach for real-time control, offering flexibility in handling arbitrary dynamics and cost functions. However, the original MPPI suffers from high-frequency noise in the sampled control trajectories, leading to actuator wear and inefficient exploration. In this work, we introduce Low-Pass Model Predictive Path Integral Control (LP-MPPI), which integrates low-pass filtering into the sampling process to eliminate detrimental high-frequency components and improve the effectiveness of the control trajectories exploration. Unlike prior approaches, LP-MPPI provides direct and interpretable control over the frequency spectrum of sampled trajectories, enhancing sampling efficiency and control smoothness. Through extensive evaluations in Gymnasium environments, simulated quadruped locomotion, and real-world F1TENTH autonomous racing, we demonstrate that LP-MPPI consistently outperforms state-of-the-art MPPI variants, achieving significant performance improvements while reducing control signal chattering.
@@ I INTRODUCTION
In general, there are two main approaches to MPC: sampling-based and optimization-based. Optimization-based MPC algorithms provide an efficient way to find optimal control sequences using dynamics and cost function gradients [5, 6]. However, they typically impose requirements on the dynamics model or cost function formulations, such as differentiability or continuity. An interesting alternative is the sampling-based MPC. One of the main benefits of this approach is that the dynamics and cost functions can be arbitrary, and the only requirement is to evaluate them relatively fast. The two algorithms within this group, which have proven their effectiveness in numerous tasks, such as off-road driving [2], drone flight [7], and control of high-dimensional simulated systems (humanoids, dexterous hands, manipulators) [8, 9], are the Cross Entropy Method (CEM) [8, 10] and Model Predictive Path Integral Control (MPPI) [2].
The core of both MPPI and CEM approaches is to evaluate the performance of the control trajectories sampled from a sequence of Gaussian distributions. This approach results in the lack of temporal correlation within the individual trajectories (see the red part of Figure 1). In fact, these trajectories are approximately the white noise signals, i.e., they are evenly composed of all possible frequencies. This, in turn, results in the overrepresentation of the high-frequency components when compared to the expected optimal behaviors in most of the robotic systems. Moreover, the dynamics of most robots dampen these components significantly, so they do not significantly affect the obtained costs. Thus, they do not contribute to the efficient exploration of the control signals domain but instead result in chattering of the applied controls and wear out of the actuators.
Figure 1: The proposed Low-Pass Model Predictive Path Integral (LP-MPPI) control smooths out the controls via the low-pass filtering and increases the efficiency of the MPPI.
@@
In this paper, we propose to address these issues by extending the MPPI algorithm with a low-pass filter acting on the sampled control trajectories. Our method, called Low-Pass Model Predictive Path Integral control (LP-MPPI), is presented schematically in Figure 1. Our approach cuts off the harmful and ineffective high-frequency components of the control signal perturbations, increasing the efficiency of the MPPI algorithm and smoothing the resultant control signal. Moreover, unlike colored noise, it does not bias the control signal frequencies in the filter passband, allowing for efficient exploration of performant trajectories in the assumed range of admissible frequencies. Our method introduces only two interpretable parameters with a clear physical meaning, allowing intuitive tuning. They directly control the frequency range in which the search for the optimal control trajectory is focused and the damping characteristics beyond this range. Finally, our method is easy to implement and does not introduce significant computational overhead compared to the MPPI algorithm.
Due to space limitations, this paper focuses on the family of MPPI algorithms. However, given the structural similarities of sampling-based MPC we suppose that our filtering approach can be successfully applied to other algorithms, e.g. CEM.
We perform an extensive experimental evaluation of the proposed method against the State-of-the-Art variants of the MPPI in three Gymnasium environments over a wide range of the MPPI parameters. The results show that our approach consistently outperforms the considered baselines, improving their results by 24% on average. Moreover, it reduces the chattering in the control signals, obtaining the smallest difference between the computed control signals and those filtered with a Savitzky-Golay filter among the considered methods. Furthermore, in our second experiment, we assess the ability of the proposed method to control more complex systems, such as quadruped robots, by combining our LP-MPPI algorithm with the recently developed Dial-MPC [14], showing how easily and effectively the proposed approach can be integrated with the other MPPI-based methods. In this case, LP-Dial-MPC outperformed Dial-MPC by more than 32% on average. Finally, we evaluate our approach in real-world autonomous F1TENTH racing, outperforming all of the baselines, most of them by a large margin.
The contributions of this paper can be summarized as follows:
  • We propose a Low-Pass Model Predictive Path Integral algorithm, which enables shaping the frequency spectrum of the control trajectories distribution, reducing the impact of high-frequency noise, while maintaining the capability of effective exploration within the assumed frequency range.
  • We conduct a thorough experimental analysis of the proposed approach in several simulated environments, showing that the proposed method consistently outperforms the State-of-the-Art MPPI-based control approaches, while reducing the amount of high-frequency components in the applied control signals.
@@ II Related Work
In recent years, many attempts have been made to improve the effectiveness of the MPPI algorithm [9, 11, 12, 13, 14]. A substantial number of these methods were considering how to increase the smoothness of the control signal and improve the sampling exploration efficiency.
In [12], authors proposed to smooth the control signals generated by MPPI, by (i) lifting the input-space, i.e., sampling in the space of the derivatives of the actual controls, and (ii) adding a smoothness cost. However, lifting the input space can decrease the exploration efficiency, especially for small $dt$. Our frequency-based shaping of control signal perturbations can adjust the time correlation of the trajectories independently of $dt$ and does not require additional cost terms. Moreover, the authors of [12] showed that using a low-pass filter on weighted perturbations before updating the nominal control trajectory results in a poorer performance than their proposed SMPPI. Our work shows that if the low-pass filtering is applied before the control perturbations are simulated, it outperforms SMPPI.
@@
Another interesting strategy to smooth the control signals and improve exploration efficiency is to exploit spline interpolation [9, 11, 15]. Instead of sampling a long sequence of actions, one can sample only a reduced number of spline control points and then interpolate between them. Thus, the dimension of the exploration space is reduced and the resulting trajectories are smoother. However, in this approach, the control over the smoothness is indirect and rather coarse and may require predicting the time increments along the controls [15].
@@
A more direct approach to improving control smoothness and exploration capabilities is to exploit some specific distributions from which the control perturbation samples are drawn, as the default Gaussian one may not be an optimal choice. In this theme, the authors of [16] propose to use a mixture of the original normal distribution with the log-normal distribution to increase the probability of sampling the larger deviations from the nominal control sequence to improve the exploration. In turn, in [17] and [18] authors explored the use of multimodal distributions, i.e., Gaussian Mixture Models and empirical particle-based distributions with Stein Variational Gradient Descent updates, to increase the exploration capabilities by allowing one to maintain multiple hypotheses. Recently, learning-based approaches have also been used to determine better sampling distributions based on the structure of the environment [19, 20], enabling exploration focused on promising areas of the state space. Nevertheless, none of these methods considers the correlation between subsequent controls within each sampled control trajectory perturbation, which may result in a chattering of the control signal.
@@
Methods that adjust the sampling distribution to address this particular issue, while increasing the exploration capabilities of the sampling-based MPC, were proposed in [8] and [13]. The approach presented in [8], called iCEM, is an improved version of the Cross-Entropy Method (CEM) [10]. It shows that sampling the actions from the colored-noise distributions significantly improves the original CEM approach. The authors of [13] employed the same idea in the MPPI algorithm and have shown improved smoothness and performance. In our work, we also focus on shaping the MPPI sampling distribution in the frequency domain; however, instead of using colored noise distributions, we use a low-pass filter that gives the user more degrees of freedom and allows uniform sampling within the desired range of low frequencies.
@@ III-A Problem definition
In this paper, we consider an optimal control problem of the form
$\min_{u_{t:t+H}}$ $J(x_{t:t+H+1},u_{t:t+H})=\sum_{h=0}^{H}c(x_{t+h},u_{t+h})+c_{f}(x_{t+H+1})$
subject to (1)
$x_{t+h+1}=f(x_{t+h},u_{t+h}),\quad\forall h\in\{0,...,H\}$
$x_{t:t+H+1}\in\mathcal{X},u_{t:t+H}\in\mathcal{U},$
where $x_{t}$ and $u_{t}$ are the state and control signal at time $t$, $f$ is the system dynamics, $c$ and $c_{f}$ are step and terminal cost functions, $J$ is the total cost of the trajectory, while $\mathcal{X}$ and $\mathcal{U}$ are the sets of admissible states and controls, respectively.
@@ III-B Model Predictive Path Integral Control
Our proposed method, Low-Pass Model Predictive Path Integral Control (LP-MPPI), is based strongly on the original MPPI algorithm [2]. Therefore, we recall its pseudocode in Algorithm 1. In general, the idea is to (i) draw multiple control trajectories around the current nominal control trajectory, (ii) simulate them using the model of the system, (iii) compute their costs, and finally (iv) update the nominal trajectory based on the costs obtained by the perturbed controls. In this paper, we focus on the commonly overlooked aspect of the MPPI algorithm – drawing random control samples. In fact, the only part of the algorithm that we would like to analyze and improve is located in line 4. To do so, we will analyze the original MPPI algorithm [2] and its recent extension [13] from a frequency domain perspective.
@@ III-C Spectral analysis
In the original MPPI, all elements of the noise sequences are sampled from the Gaussian distribution with some constant covariance matrix $\Sigma$. This results in a lack of correlation between the subsequent elements of the noise sequence. In general, this may be seen as an advantage, as perturbing the system with white noise is a well-known technique in system identification [21], which ensures that the system dynamics is excited with all possible control frequencies. In the context of MPC, similar implications should also be true for the cost function $J$, which depends on both controls and state trajectories. However, in the context of robotic applications, it is pretty uncommon to control systems that require an excitation with high-frequency noise, as most robots filter it out. Therefore, control trajectories with high-frequency noise and low-frequency noise result in similar costs, which may result in frequent changes of the following actions, and thus jittering in the real system. Moreover, we expect that for most robotic systems, the spectrum of frequencies of the optimal control trajectories is not uniformly distributed, so the use of white noise may result in a performance decrease. In fact, using the uncorrelated noise distributions corresponds to the maximum exploration in the frequency spectrum.
To limit the amount of high-frequency noise in the sampled controls and to address the lack of exploitation in the original MPPI, the authors of [13] proposed biasing the spectrum of the control trajectory perturbations towards low frequencies. As a result, they obtained an exploitation strategy that puts most of the signal energy in the lowest possible frequencies and gradually reduces the energy of the higher-frequency components. In turn, in this paper, we would like to analyze an alternative approach – low-pass filtered noise, which balances exploration and exploitation and can be controlled by an adjustable parameter – cutoff frequency.
@@
To analyze each of the aforementioned control sampling strategies and provide the rationale for the approach proposed in this paper, we conducted a small experiment. We compared the power density spectra of (i) white noise, (ii) colored noise, and (iii) low-pass filtered white noise, to the spectrum of the control signals generated by a trained RL policy from the StableBaselines3-zoo library [22] for two sample environments, Ant-v3 and Humanoid-v3. We chose the RL agent controls spectrum as a reference, as it may be considered to be very close to the frequency statistics of the optimal behaviors in these environments. To present the results in a compact form, we averaged the spectra of the individual joints for the RL agent. In addition, we averaged the spectra over 100 seeds to reduce noise. Moreover, to ensure a fair comparison between the analyzed sampling distributions, we optimized their parameters to minimize the norm between their and the RL agent action spectra. The results of this comparison are presented in Figure 2. One can see that in the considered locomotion tasks, the RL agent does not perform high-frequency actions; however, following only the lowest frequencies is also far from optimal, as it ignores the frequency bumps that occur around $2\text{\,}\mathrm{Hz}$. In fact, the spectrum that best matches the RL agent is the low-pass-filtered white noise, as it does not damp the signal at the most important frequencies. Moreover, what is also important in the context of MPPI, it allows for efficient exploration in the assumed bandwidth.
@@ III-D Low-pass Model Predictive Path Integral Control
Following the observations made in the previous section, we introduce the Low-pass Model Predictive Path Integral Control (LP-MPPI). The core of the proposed algorithm is the introduction of a temporal correlation between the subsequent control perturbations by using a low-pass filter on them. The proposed approach is formalized in Algorithm 2, where the changes with respect to the original MPPI algorithm are marked in blue. Our proposed method introduces a small yet important modification, which, first, biases the search of the best control trajectories to the low-frequency signals and, second, gives the algorithm user the capability to control the trade-off between the exploration and exploitation in the frequency domain. These two phenomena can be controlled by tuning the parameters of the low-pass filter, which we describe in detail in Section III-E. Importantly, the proposed modifications do not introduce significant computational overhead to the MPPI algorithm, as filtering typically requires a small number of multiplications proportional to the length of the horizon $H$ and, in most cases, is significantly cheaper than evaluating the system’s dynamics.
In our software implementation, we utilized a simple and computationally efficient digital implementation of the Butterworth filter [23]. This type of filter is maximally flat in the passband, which prevents introducing biases in terms of the specific frequencies within the desired range. In turn, its frequency spectrum is not as steep as for example for Chebyshev filters, however, as can be seen in Figure 2, this property is not necessarily needed in the context of sampling performant control trajectories. Similarly, the relatively high phase shift introduced by the Butterworth filter does not pose a significant issue in our particular application, as we filter the white noise signals (shifting temporally uncorrelated signals does not change anything).

### v3

# LP-MPPI: Low-Pass Filtering for Efficient Model Predictive Path Integral Control

@@ Abstract
Model Predictive Path Integral (MPPI) control is a widely used sampling-based approach for real-time control, valued for its flexibility in handling arbitrary dynamics and cost functions. However, it often suffers from high-frequency noise in the sampled control trajectories, which hinders the search for optimal controls and transfers to the applied controls, leading to actuator wear. In this work, we introduce Low-Pass Model Predictive Path Integral Control (LP-MPPI), which integrates low-pass filtering into the sampling process to eliminate detrimental high-frequency components and enhance the algorithm’s efficiency. Unlike prior approaches, LP-MPPI provides direct and interpretable control over the frequency spectrum of sampled control trajectory perturbations, leading to more efficient sampling and smoother control. Through extensive evaluations in Gymnasium environments, simulated quadruped locomotion, and real-world F1TENTH autonomous racing, we demonstrate that LP-MPPI consistently outperforms state-of-the-art MPPI variants, achieving significant performance improvements while reducing control signal chattering.
@@ I INTRODUCTION
In general, there are two main approaches to MPC: sampling-based and optimization-based. Optimization-based MPC algorithms provide an efficient way to find optimal control sequences using dynamics and cost function gradients [5, 23]. However, they typically impose requirements on the dynamics model or cost function formulations, such as differentiability or continuity. An interesting alternative is sampling-based MPC. One of the main benefits of this approach is that the dynamics and cost functions can be arbitrary, and the only requirement is to evaluate them relatively fast. The two algorithms within this group, which have proven their effectiveness in numerous tasks, such as off-road driving [27], drone flight [14], and control of high-dimensional simulated systems (humanoids, dexterous hands, manipulators) [17, 2], are the Cross Entropy Method (CEM) [20] and Model Predictive Path Integral Control (MPPI) [27].
The core of both MPPI and CEM approaches is to evaluate the performance of the control trajectories sampled from a sequence of Gaussian distributions. This approach results in a potential lack of temporal correlation within individual trajectories (see the red part of Figure 1). In fact, control trajectory perturbations are approximately white noise signals, i.e., they are evenly composed of all possible frequencies. This, in turn, results in the overrepresentation of the high-frequency components when compared to the expected optimal behaviors in most robotic systems. Moreover, the dynamics of most robots naturally dampen these components, so their impact on the resulting performance is minimal. Thus, they do not contribute to the efficient search for optimal control signals but instead result in chattering of the applied controls and wear out of the actuators.
Figure 1: The proposed Low-Pass Model Predictive Path Integral (LP-MPPI) control smooths out the controls via the low-pass filtering of control perturbations and increases the efficiency of the MPPI.
@@
In this paper, we address these issues by extending the MPPI algorithm with a low-pass filter applied to the sampled control perturbations. Our method, called Low-Pass Model Predictive Path Integral control (LP-MPPI), is presented schematically in Figure 1. Our approach eliminates the harmful and ineffective high-frequency components of the control signal perturbations, increasing the efficiency of the MPPI algorithm and smoothing the resultant control signal without sacrificing its responsiveness. Moreover, unlike colored noise, it does not bias the control signal frequencies in the filter’s passband, enabling an effective search for high-performing trajectories across the admissible frequency range. Our method introduces only two interpretable parameters with clear physical meaning, allowing intuitive tuning. They directly control the frequency range in which the search for the optimal control trajectory is focused and the damping characteristics beyond it. Finally, LP-MPPI is easy to implement and adds negligible computational overhead compared to the MPPI algorithm.
We conduct an extensive experimental evaluation of the proposed method against the State-of-the-Art variants of the MPPI in three Gymnasium environments, over a wide range of MPPI parameters. The results demonstrate that our approach consistently outperforms the considered baselines, improving their results by 24% on average, while significantly reducing chattering in the applied control signals. To further assess its applicability, we integrate LP-MPPI with the recently developed Dial-MPC [28] and evaluate the resulting LP-Dial-MPC on quadruped robots. This experiment highlights the ease and effectiveness with which our method can be combined with other MPPI-based approaches, yielding an average performance gain of over 32% compared to Dial-MPC. Finally, we validate our approach in real-world autonomous F1TENTH racing, where it outperforms all baselines, most of them by a large margin.
Our contributions can be summarized as follows:
  • We propose a Low-Pass Model Predictive Path Integral algorithm, which enables shaping the frequency spectrum of the control trajectories’ perturbations distribution, improving search efficiency and reducing the high-frequency noise in the applied control signal.
  • We conduct a thorough experimental analysis of the proposed approach in several simulated environments, showing that the proposed method consistently outperforms the state-of-the-art MPPI-based control approaches, while reducing the amount of high-frequency components in the applied control signals (see Tab. I).
@@ II Related Work
In recent years, many attempts have been made to improve the effectiveness of the MPPI algorithm [2, 15, 11, 24, 28]. A substantial number of these methods considered how to increase the smoothness of the control signal and improve the sampling efficiency.
In [11], authors proposed to smooth the control signals generated by MPPI, by (i) lifting the input-space, i.e., sampling in the space of the derivatives of the actual controls, and (ii) adding a smoothness cost. This approach effectively transforms the frequency spectrum of the control signals’ perturbations into colored noise and gives only minimal control of the frequency spectrum. Our frequency-based shaping of control signal perturbations can adjust the time correlation of the trajectories in a more fine-grained way, does not introduce such a strong bias to the lowest frequencies, and does not require additional cost terms. Moreover, it should be noted that adding smoothness costs reduces search efficiency because part of the drawn trajectories will be effectively excluded due to high smoothness costs.
Interestingly, the authors of [11] showed that using a low-pass Savitzky-Golay filter on the control trajectory update signal (see MPPI(SGF) in [11]) results in a poorer performance than their proposed SMPPI. Our work shows that if the low-pass filtering is applied before the control perturbations are simulated, it maintains constraint satisfaction, introduces no phase distortion, and outperforms SMPPI [11].
@@
Another interesting strategy to smooth the control signals and improve exploration efficiency is to exploit spline interpolation [2, 15, 6]. Instead of sampling a long sequence of actions, one can sample only a reduced number of spline control points and then interpolate between them. Thus, the dimension of the exploration space is reduced, and the resulting trajectories are smoother. However, in this approach, the control over the smoothness is indirect and rather coarse and may require predicting the time increments along the controls [6, 10].
@@
A more direct approach to improving control smoothness and exploration capabilities is to exploit some specific distributions from which the control perturbation samples are drawn, as the default Gaussian one may not be an optimal choice. In this theme, the authors of [16] propose to use a mixture of the original normal distribution with the log-normal distribution to increase the probability of sampling the larger deviations from the nominal control sequence to improve the exploration. In turn, in [25] and [13], authors explored the use of multimodal distributions, i.e., Gaussian Mixture Models and empirical particle-based distributions with Stein Variational Gradient Descent updates, to increase the exploration capabilities by allowing one to maintain multiple hypotheses. Recently, learning-based approaches have also been used to determine better sampling distributions based on the structure of the environment [18, 21], enabling focusing the search on promising areas of the state space. Nevertheless, none of these methods considers the correlation between subsequent controls within each sampled control trajectory perturbation, which may result in a chattering of the control signal.
@@
Methods that adjust the sampling distribution to address this particular issue, while increasing the exploration capabilities of the sampling-based MPC, were proposed in [17] and [24]. The approach presented in [17], called iCEM, is an improved version of the Cross-Entropy Method (CEM) [20]. It shows that sampling the actions from the colored-noise distributions significantly improves the original CEM approach. The authors of [24] employed the same idea in the MPPI algorithm, demonstrating improved smoothness and performance. In our work, we also focus on shaping the MPPI sampling distribution in the frequency domain. However, instead of relying on colored noise, we use a low-pass filter, which provides greater flexibility and enables uniform sampling across the desired range of low frequencies.
@@ III-A Problem definition
We consider the following optimal control problem $\min_{u_{t:t+H}}$ $J(x_{t:t+H+1},u_{t:t+H})=\sum_{h=0}^{H}c(x_{t+h},u_{t+h})+c_{f}(x_{t+H+1})$ subject to (1) $x_{t+h+1}=f(x_{t+h},u_{t+h}),\quad\forall h\in\{0,...,H\}$ $x_{t:t+H+1}\in\mathcal{X},u_{t:t+H}\in\mathcal{U},$ where $x_{t}$ and $u_{t}$ are the state and control signal at time $t$, $f$ is the system dynamics, $c$ and $c_{f}$ are step and terminal cost functions, $J$ is the total cost of the trajectory, while $\mathcal{X}$ and $\mathcal{U}$ are the sets of admissible states and controls, respectively.
@@ III-B Model Predictive Path Integral Control
Our proposed method, Low-Pass Model Predictive Path Integral Control (LP-MPPI), is based strongly on the original MPPI algorithm [27]. Therefore, we recall its pseudocode in Algorithm 1. In general, the idea is to (i) draw multiple control trajectories around the current nominal control trajectory, (ii) simulate them using the model of the system, (iii) compute their costs, and finally (iv) update the nominal trajectory based on the costs obtained by the perturbed controls. In this paper, we focus on the commonly overlooked aspect of the MPPI algorithm – drawing random control perturbations. In fact, the only part of the algorithm that we would like to analyze and improve is located in line 4. To do so, we will analyze the original MPPI algorithm [27] and its recent extension [24] from a frequency domain perspective.
@@ III-C Spectral analysis
In the original MPPI, all elements of the noise sequences are sampled from the Gaussian distribution with some constant covariance matrix $\Sigma$. This results in a lack of correlation between the subsequent elements of the noise sequence. In general, this may be seen as an advantage, as perturbing the system with white noise is a well-known technique in system identification [9], which ensures that the system dynamics are excited with all possible control frequencies. In the context of MPC, similar implications should also be true for the cost function $J$, which depends on both controls and state trajectories. However, in the context of robotic applications, it is pretty uncommon to control systems that require an excitation with high-frequency noise, as most robots filter it out. Therefore, control trajectories with high-frequency noise and low-frequency noise result in similar costs, which may result in frequent changes of the following actions, and thus jittering in the real system. Moreover, we expect that for most robotic systems, the spectrum of frequencies of the optimal control trajectories is not uniformly distributed, so the use of white noise may reduce the search efficiency and result in a performance decrease. In fact, using the uncorrelated noise distributions corresponds to the maximum exploration in the frequency spectrum.
To limit the amount of high-frequency noise in the sampled controls, the authors of [24] proposed biasing the spectrum of the control trajectory perturbations towards low frequencies. As a result, they obtained a strategy that puts most of the signal energy in the lowest possible frequencies and gradually reduces the energy of the higher-frequency components. In turn, in this paper, we analyze an alternative approach based on low-pass filtered noise, which balances the bias toward lower frequencies with efficient exploration of the admissible spectrum, and enables explicit control of this trade-off through adjustable parameters – cutoff frequency and filter order.
@@
To analyze each of the aforementioned control sampling strategies and provide the rationale for the approach proposed in this paper, we conducted an experiment. We compared the power density spectra of (i) white noise, (ii) colored noise, and (iii) low-pass filtered white noise, to the spectrum of the control signals generated by a trained RL policy from the StableBaselines3-zoo library [19] for two sample environments, Ant-v3 and Humanoid-v3. We chose the RL agent controls spectrum as a reference, as it may be considered to be very close to the frequency statistics of the optimal behaviors in these environments. To present the results in a compact form, we averaged the spectra of the individual joints for the RL agent. In addition, we averaged the spectra over 100 seeds to reduce noise. Moreover, to ensure a fair comparison between the analyzed sampling distributions, we optimized their parameters to minimize the norm between their and the RL agent action spectra. The results of this comparison are presented in Figure 2. One can see that in the considered locomotion tasks, the RL agent does not perform high-frequency actions; however, following only the lowest frequencies is also far from optimal, as it ignores the frequency bumps that occur around $2\text{\,}\mathrm{Hz}$. In fact, the spectrum that best matches the RL agent is the low-pass-filtered white noise, as it does not damp the signal at the most important frequencies. Moreover, what is also important in the context of MPPI, it allows for efficient exploration in the assumed bandwidth.
@@ III-D Low-pass Model Predictive Path Integral Control
Following the observations made in the previous section, we introduce the Low-pass Model Predictive Path Integral Control (LP-MPPI). The core of the proposed algorithm is the introduction of a temporal correlation between the subsequent control perturbations by using a low-pass filter on them. The proposed approach is formalized in Algorithm 2, where the changes with respect to the original MPPI algorithm are marked in blue. Our method introduces a subtle yet important modification that (i) biases the search for optimal control trajectories toward low-frequency signals and (ii) allows the user to control the frequency spectrum within which the algorithm searches for candidate trajectory updates. These features can be precisely controlled by tuning the parameters of the low-pass filter, which we describe in detail in Section III-E. Importantly, the proposed modifications do not introduce significant computational overhead to the MPPI algorithm, as filtering typically requires a small number of multiplications proportional to the horizon length $H$ and, in most cases, is significantly cheaper than evaluating the system’s dynamics.
One of the crucial design choices is to decide which signals should be smoothed in order to improve efficiency and reduce chattering without sacrificing controller reactiveness or violating system constraints. In the proposed approach, we apply low-pass filtering to sampled noise sequences, i.e. control trajectory perturbations (see line 2 of Algorithm 2). This way, we strongly bias the search for control trajectories towards low frequencies, but we do not constrain the applied control signal frequency spectrum. The first control may change rapidly, but the change of the entire control trajectory is correlated. Thus, the search is biased, but the applied control signal maintains its reactiveness. Moreover, unlike existing filtering approaches, such as MPPI(SGF) [11], our approach improves search efficiency by evaluating only smoothed constraint-compliant perturbations and does not introduce any phase shift to the applied control signal, as if the filtering were applied directly to it.
In our software implementation, we utilized a simple and computationally efficient digital implementation of the Butterworth filter [3]. This type of filter is maximally flat in the passband, which prevents introducing biases to specific frequencies within the desired range. In turn, its frequency spectrum is not as steep as, for example, that of Chebyshev filters; however, as can be seen in Figure 2, this property is not necessarily needed in the context of sampling performant control trajectories. Similarly, the relatively high phase shift introduced by the Butterworth filter is not problematic in our approach, since temporally uncorrelated perturbations remain unaffected by such shifts.